OpenStudy (anonymous):

Integrate (cosxsinx) / (cos^2(x) - 1)

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ cosxsinx }{ \cos^2x -1 }dx\]

hartnn (hartnn):

u can write cos^2 x as 1-sin^2 x

OpenStudy (anonymous):

ok\[\frac{\sin x \cos x}{\cos^2 x-1}=\frac{1}{2}\frac{\sin 2x }{\frac{1+\cos 2x}{2}-1}\]

hartnn (hartnn):

maybe no need to go for 2x angle.

OpenStudy (lgbasallote):

let u = cos^2 x - 1 du = 2cosx sinx it's much easier...

OpenStudy (anonymous):

sinxcosx/(-sin^2x)

OpenStudy (anonymous):

-cotx

OpenStudy (anonymous):

lol...mine is much harder

OpenStudy (lgbasallote):

\[\implies \frac 12 \int \frac 1u du\]

hartnn (hartnn):

then u can cancel sin x from numerator and denominator to get -cot x

OpenStudy (lgbasallote):

i'll organize my solution... \[\int \frac{\cos x \sin x}{\cos^2 x - 1}dx\] let \[u = \cos^2 x - 1\] \[du = -2\sin x \cos x dx\] so the integral becomes \[\implies -\frac 12 \int \frac {du}u\]

OpenStudy (lgbasallote):

integrate that

OpenStudy (anonymous):

let (cosx)^2 = t 2(cosx)(-sinx)dx = dt so given eq is integral[1/2(1/(1-t^2)]dt

OpenStudy (anonymous):



OpenStudy (lgbasallote):

if you let cos^2 x = t then your denominator should be t - 1 @LeeYeongKyu

OpenStudy (anonymous):

no - is calculated in dominater.

OpenStudy (lgbasallote):

or if you factor out negative...should be 1-t

OpenStudy (anonymous):

(cosx)^2 = -2cosxsinx

OpenStudy (anonymous):

if you differentite it

OpenStudy (lgbasallote):

you let cos^2 x = t so your denominator should be 1 - t

OpenStudy (anonymous):

sorry... i miss wrote.. thanx

OpenStudy (anonymous):

@LeeYeongKyu http://i.imgur.com/AaCa1.gif

OpenStudy (anonymous):

i had a little hasitate

OpenStudy (lgbasallote):

here's another solution.. \[\int \frac{\cos x \sin x}{\cos^2 x - 1}dx\] let \[u = \cos x\] \[du = -\sin xdx\] so integral becomes.. \[\implies -\int \frac{udu}{u^2 - 1}\] it's harder..bot possible

OpenStudy (lgbasallote):

but*

OpenStudy (anonymous):

no it's easy thing to integrate...

OpenStudy (lgbasallote):

i meant harder than what i first wrote

OpenStudy (anonymous):

right

OpenStudy (anonymous):

the answer I've been getting is \[\frac{ 1 }{ 2}\ln \left| \cos^2x-1\right| + c\], the correct answer according to the book is \[\ln \left|cscx\right| + c\]

OpenStudy (lgbasallote):

\[\frac 12 \ln | \cos^2x - 1| + c \implies \frac 12 \ln |-\sin^2 x| + c\] \[\implies \frac 12 \ln | \csc^2 x| + c\] \[\implies \ln |\csc x| + c\] did you follow that?

OpenStudy (anonymous):

so \[\cos^2x-1 = -\sin^2x = \csc^2x\]?

OpenStudy (lgbasallote):

nevermind that..

OpenStudy (lgbasallote):

\[\cos^2 x - 1 = -(1-\cos^2 x) = -\sin ^2 x\] right?

OpenStudy (anonymous):

I'm really bad at my trig identities, if you say so I'll believe it

OpenStudy (lgbasallote):

this came from \[\sin^2 x + \cos ^2 x = 1\] familiar?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok I see it now

OpenStudy (lgbasallote):

wonderful

OpenStudy (lgbasallote):

so you have \[\frac 12 \ln | -\sin^2 x| + c\] right?

OpenStudy (anonymous):

right

OpenStudy (lgbasallote):

this is why i told you that you misplaced a negative sign... in this case..the absolute value will just cancel out the negative sign...so you can't get csc^2 x

OpenStudy (lgbasallote):

your integral should have been \[-\frac 12 \ln | \cos^2 x - 1| + c\] because this gives you \[-\frac 12 \ln | \sin ^2 x| + c\] if you use power rule here... \[\implies \ln |(\sin ^2 x)^{-1/2} | + c\] that becomes \(\ln |\csc x| + c\] does that make sense to you?

OpenStudy (lgbasallote):

or should i try a different approach to make you see what i mean?

OpenStudy (anonymous):

ok I see, you used clnx = lnx^c so that we could get 1/sin^2x which equals cscx

OpenStudy (lgbasallote):

right!

OpenStudy (anonymous):

:D thanks so much for sticking it out to the end!

OpenStudy (lgbasallote):

welcome